OpenStudy (anonymous):

The formula for the circumference of a circle is C = 2r, where r is the radius and C is the circumference. The equation solved for r is r = . Find the radius of a circle that has a circumference of 16.

OpenStudy (rizags):

thats the wrong formula for circumference

OpenStudy (jhannybean):

\[C=2\pi r\]

OpenStudy (rizags):

circumference is C=2πr

OpenStudy (rizags):

2(pi)r

OpenStudy (jhannybean):

So if you are solving for r, then \[r = \frac{C}{2\pi}\]

OpenStudy (jhannybean):

\[C = 16 ~,~ r = \frac{C}{2\pi} \implies r = ~?\]